# Boosting Scientific Error-Bounded Lossy Compression through Optimized Synergistic Lossy-Lossless Orchestration

### v1

@@ Abstract.
copyright: none
@@ 1. Introduction
No doubt that we are in the exascale computing era, in which today’s scientific applications tend to be running on extremely large-scale environments, continuously producing masses of scientific data. The world’s best exascale supercomputers, like Frontier (fro, 2022), Aurora (aur, 2024), and El Capitan (elc, 2023), which present peak performances around or over 1 Exaflops, have pushed the scientific computational power to a brand new magnitude, which is highly attributed to the cutting-edge GPU hardware that those supercomputers facilitate. Correspondingly, exascale scientific computing workflows have been established on GPU platforms, which include diverse scientific domains such as cosmology (NYX simulation, 2019), turbulence (Yeung et al., 2025), and molecular dynamics (Carrasco-Busturia et al., 2024). However, higher computational power always leads to higher data generation speed, and the development of scientific computing workflows on GPU platforms is in great need of more effective data management solutions, which are critical to the storage, analysis, and transmission of exascale scientific data. For example, in (Yeung et al., 2025), a GPU-accelerated turbulence simulation framework is proposed, which can generate data grids of 35 trillion grid points (128 TB in single-precision). To this end, effective scientific data reduction techniques are required to be available on GPU platforms to relieve the system bottleneck in I/O and storage with relatively low latency. The error-bounded lossy compression of scientific data can greatly reduce the data volume while preserving the per-data accuracy well, so it is commonly regarded as the best-fit data reduction solution for scientific data (Liang et al., 2023; Lindstrom, 2014; Tian et al., 2020; Zhao et al., 2020a; Tao et al., 2019).
@@
Over the past years, there have been various practices and deliverables of GPU-based scientific error-bounded lossy compression, which project substantial potential in integrating them into GPU-based scientific computing workflows. For example, cuSZ (Tian et al., 2020, 2021; Liu et al., 2024c) and cuZFP (Lindstrom, 2014; cuZFP, 2020) have been leveraged by existing scientific applications, and cuSZp (Huang et al., 2023, 2024a) is an excellent high-throughput data compressor to be leveraged in performance-favored GPU data compression. Nevertheless, a significant research gap remains in this field: An open-source design for high-ratio scientific error-bounded lossy compression on GPUs is still missing. Most existing works focus on high throughput, relying on low-cost and low-quality data processing algorithms. cuSZ-I (Liu et al., 2024c), as the first and the only existing GPU-based high-ratio-and-quality scientific data compressor, integrates an NVIDIA-proprietary lossless encoding module to achieve a compression ratio higher than 32, yet still exhibits much lower compression ratio than CPU-based compressors, like SZ3 (Liang et al., 2023) and QoZ (Liang et al., 2023; Liu et al., 2024a). The most critical challenges of high-ratio scientific error-bounded lossy compression on GPU platforms are: 1) It is non-trivial to propose effective GPU-parallelization models and practical high-throughput implementations of high-ratio scientific data compression algorithms; 2) To optimize the GPU-based high-ratio scientific lossy compression pipelines, a thorough and in-depth benchmarking and analysis on the data decomposition and encoding modules on GPU platforms are required.
@@
In this paper, committing to optimizing high-ratio scientific error-bounded lossy compression on GPU platforms, we propose cuSZ-Hi, which is a general, open-source, and highly effective solution. From the decomposition-encoding perspective of error-bounded lossy compression, based on the prior success of interpolation-based error-bounded lossy compressors (Zhao et al., 2021; Liu et al., 2022; Liu et al., 2024a, c), we fine-craft a well-rounded GPU-parallelized interpolation-based data predictor, which is oriented from maximizing the lossless compressibility of integer decomposition data. Furthermore, from systematic investigation, evaluation, and benchmarking on GPU-based numerical lossless encoders, we establish the best lossless pipelines for cuSZ-Hi, which fully leverages different aspects of redundancy information in the data, ensuring the high-magnitude compression ratio among all use cases. Our contributions are as follows:
@@
  • We perform a comprehensive evaluation of cuSZ-Hi on 6 real-world scientific datasets. cuSZ-Hi exhibits outperforming compression ratio and quality compared to all existing state-of-the-arts. cuSZ-Hi can achieve up to 249% compression ratio improvement over existing scientific lossy compressors under the same error bound and up to 215% compression ratio improvement under the same decompression data PSNR.
We organize the rest of this paper as follows: Section 2 discusses the related work of this paper. Section 3 formulates our research target and the background of interpolation-based scientific error-bounded lossy compression on GPUs. In Section 4, we propose the framework design and module composition of cuSZ-Hi. Section 5 demonstrates our detailed designs for optimizing cuSZ-Hi. In Section 6, we present the evaluations for cuSZ-Hi. Section 7 concludes this work and discusses future plans.
@@ 2.1. Scientific error-bounded lossy compression
Error-bounded lossy compression has been regarded as the best-fit strategy for scientific data size reduction. Existing error-bounded lossy compressors have proposed diverse technical solutions, and present different capabilities on use cases and metrics. Among the most representative works, SZ-family (Zhao et al., 2020b; Liang et al., 2023; Liu et al., 2022; Liu et al., 2024a) follows a hybrid design of spline interpolation (Zhao et al., 2021), Linear Regression, and Lorenzo extrapolation (Tao et al., 2017). ZFP (Lindstrom, 2014) and SPERR (Li et al., 2023) proceed data with transforms and specialized encoding schemes. TTHRESH (Ballester-Ripoll et al., 2019) leverages high-order singular value decomposition to effectively de-correlate the input data. To maximally reduce the compressed data size, there are various deep-learning-integrated frameworks for scientific error-bounded lossy compression (Liu et al., 2021a, 2023; Huang and Hoefler, 2022; Lu et al., 2021; Hayne et al., 2021; Liu et al., 2024b; Li et al., 2024). To optimize the compression throughput, compressors like SZx (Yu et al., 2022) and SZp (FZ-light) (Huang et al., 2024b) reduce their lossy data processing scheme to very-simplified offset-computation and bit-analysis.
@@ 2.2. Error-bounded lossy compression on GPUs
  • cuSZ: cuSZ (Tian et al., 2020, 2021; Liu et al., 2024c) was proposed in 2020, revised in 2021, and updated in 2024. The latest version of cuSZ merges the recently proposed cuSZ-I (Liu et al., 2024c) and the original Lorenzo-based cuSZ (Tian et al., 2020, 2021). It integrates 2 data predictors (Lorenzo predictor and interpolation predictor), then applies Huffman encoding (and also an optional NVIDIA Bitcomp) as the lossless module.
  • cuSZp2: cuSZp2 (Huang et al., 2024a) is an end-to-end performance optimized scientific error-bounded lossy compressor on GPU platforms. It leverages 1D offset prediction and fix-length encoding for high-speed scientific data compression.
@@
  • Others: cuSZx (Yu et al., 2022) features a monolithic design to deliver extremely high throughput at the cost of lower data quality and compression ratio. MGARD-GPU (Gong et al., 2023) is another GPU compressor practice based on the MGARD compression algorithm (Ainsworth et al., 2018).
@@ 3.1. Problem formulation of scientific error-bounded lossy compression
(1) $\begin{split}&Maximize\ \frac{|X|}{|Z|}\\ s.t.\ &Z=\operatorname{Cmp}(X),X^{\prime}=\operatorname{Dec}(Z)\\ and\ &\left\|X-X^{\prime}\right\|_{\infty}=\max|x_{i}-x_{i}^{\prime}|\leq\epsilon\\ \end{split}$
In Eq. 1, the upper bound of point-wise data error $\epsilon$ is named the error bound in the compression task. We further specify this problem in 2 aspects: 1) Regarding the GPU-based error-bounded lossy compression, the whole compression pipeline should be highly parallelized on the hardware, presenting high computing throughputs (often ten or hundreds of Gigabytes of data per second in modern platforms); 2) We observe that scientific error-bounded lossy compressors mostly follow a framework design of lossy decomposition plus lossless encoding (detailed in Section 4). The input floating-point data is first lossy decomposed to integer data with high compressibility, and the integers are further lossless encoded to gain size reduction. Eq. 2 involves this framework scheme, formulating our research target. We would like to jointly optimize the integer decomposition process (Dcp) and the lossless encoding (Enc) process while preserving the required computing throughput ($\operatorname{Speed}(\operatorname{Dcp},\operatorname{Enc})$) with speed higher than a threshold $S_{t}$ (we initially set it to a comparable speed to cuSZ-I):
@@ 3.2. Interpolation-based error-bounded lossy compression on GPUs
In this subsection, we introduce the interpolation-based error-bounded lossy compression on GPUs (specifically cuSZ-I (Liu et al., 2024c)) as our research background, as cuSZ-Hi also integrates an interpolation-based data predictor. The core design of cuSZ-I is illustrated in Figure 1. To ensure parallelization, cuSZ-I partitions the input data into small chunks (e.g., $33\times 9\times 9$ for 3D data), each composed of 4 basic blocks (shown in Figure 1), and allocates each chunk to a thread block. Leveraging spline interpolations, a thread block predicts the allocated data chunk from a losslessly stored sparse anchor grid (in the stride of 8 on each dimension) and then quantizes the prediction errors to integers for the subsequent lossless size reduction process. The interpolations follow a hierarchical scheme, performing interpolation in different sequential steps, from high levels (with large interpolation strides) to low levels (with small interpolation strides) along all dimensions. The computations are fully parallelized on the threads within each interpolation level and dimension. After acquiring the quantized integer errors (a.k.a. the quantization codes), cuSZ-I applies Huffman encoding and an optional proprietary NVIDIA Bitcomp (nvc, [n.d.]) lossless compression module to finalize the size reduction.
cuSZ-I (especially its Bitcomp-integrated version) has achieved state-of-the-art compression ratio and data quality in GPU-based scientific error-bounded lossy compression. In this paper, we will keep focusing our compressor design on interpolation and integer encoding techniques due to the unique potential of interpolation in proposing high-accuracy data predictions with relatively low computational costs. Furthermore, we will address several existing critical limitations of interpolation-based scientific data compression on GPUs, including but not limited to 1) suboptimal data-partition and anchor-sampling strategy; 2) outdated design of interpolation process and its auto-optimization; 3) lacking effective and open-source lossless encoding solution for quantized errors.
@@ 4. The Framework Design of cuSZ-Hi
The compression framework design of cuSZ-Hi is proposed in Figure 2, comparing it to other existing representative GPU-based scientific error-bounded lossy compressors. Although those compressors have diverse compression framework designs, they all share a uniform scheme, which is lossy data decomposition and lossless integer encoding, as shown in Figure 2. The first module of each framework de-correlates the input floating point data, producing integers as intermediate results. Those integers have higher compressibility than the original input, contain reduced information but preserve the error bound, and are generated by different data processing techniques such as direct value quantization (in cuSZp2), Lorenzo extrapolation (in cuSZ), and spline interpolation (in cuSZ-I). Afterward, the lossless encoding modules in each compressor reduce the size of the intermediate integers with no information loss so that the final output can serve as the error-bounded compression result. We detail the framework modules of cuSZ-Hi as follows and explain why they bring an optimized design regarding compression ratio and quality.
@@
  • Lossy data decomposition: cuSZ-Hi adopts spline-interpolation-based data prediction, which features an advanced design over the predictor in cuSZ-I (Liu et al., 2024c). Spline-interpolation-based data prediction has been proven to be the best-fit data decomposition technique for scientific error-bounded lossy compression because it can achieve both high data prediction accuracy and computation speed (Zhao et al., 2021; Liu et al., 2022; Liu et al., 2024a, c). In contrast, compressors with other lossy decomposition techniques, such as cuSZ, cuSZp2, and cuZFP, suffer from a significantly lower compression ratio than cuSZ-I. Based on the high-throughput parallelized implementation of spline interpolations proposed by (Liu et al., 2024c), cuSZ-Hi further tune and optimize the GPU-based spline interpolation design to maximize the compression ratio acquired.
@@
  • Lossless integer encoding: To maximize the compression ratio, a well-rounded lossless encoding pipeline is necessary, which should be able to leverage the data correlation information on both token-level and bit-level. Unfortunately, among existing compressors, cuSZ only integrates the frequency-based Huffman encoding and cuSZp2/cuZFP merely applies fix-length bit-wise de-redundancy encodings. cuSZ-I managed to combine both Huffman encoding and de-redundancy encoder, but the NVIDIA Bitcomp module it leverages is a proprietary product with limited usability. In cuSZ-Hi, after synthetically investigating and benchmarking available lossless encoding solutions, cuSZ-Hi proposes new open-source lossless pipelines that jointly facilitate Huffman encoding and high-performance multi-stage lossless modules. Moreover, as indicated in Figure 2, cuSZ-Hi integrates 2 different lossless pipelines, which can be flexibly selected according to the user’s requirements, presenting a higher compression ratio (in cuSZ-Hi-CR mode) or throughput (in cuSZ-Hi-TP mode).

### v2

@@ Abstract.
Data Compression, Scientific Computing, Parallel Computing
copyright: noneccs: Computing methodologies Parallel algorithmsccs: Information systems Data compressionjournalyear: 2025copyright: ccconference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759798isbn: 979-8-4007-1466-5/2025/11
[…]
by
@@ 1. Introduction
There is no doubt that we are in the exascale computing era. Today, scientific applications tend to run on extremely large-scale environments, continuously producing massive amounts of scientific data. The world’s best exascale supercomputers, like Frontier (fro, 2022), Aurora (aur, 2024), and El Capitan (elc, 2023), which present peak performances around or over 1 Exaflops, have pushed the scientific computational power to a brand new magnitude, which is highly attributed to the cutting-edge GPU hardware that those supercomputers facilitate. Correspondingly, exascale scientific computing workflows have been established on GPU platforms, which include diverse scientific domains such as cosmology (Nyx simulation, 2019), turbulence (Yeung et al., 2025), and molecular dynamics (Carrasco-Busturia et al., 2024). However, higher computational power always leads to higher data generation speed, and the development of scientific computing workflows on GPU platforms is in great need of more effective data management solutions, which are critical to the storage, analysis, and transmission of exascale scientific data. For example, in (Yeung et al., 2025), a GPU-accelerated turbulence simulation framework is proposed, which can generate data grids of 35 trillion grid points (128 TB in single-precision). To this end, effective scientific data reduction techniques are required to be available on GPU platforms to alleviate the system bottleneck in I/O and storage with relatively low latency. The error-bounded lossy compression of scientific data can significantly reduce the data volume while preserving the per-data accuracy well, so it is commonly regarded as the best-fit data reduction solution for scientific data (Liang et al., 2023; Lindstrom, 2014; Tian et al., 2020; Zhao et al., 2020a; Tao et al., 2019).
@@
Over the past years, various practices and deliverables of GPU-based scientific error-bounded lossy compression have emerged, projecting substantial potential for integrating them into GPU-based scientific computing workflows. For example, cuSZ (Tian et al., 2020, 2021; Liu et al., 2024c) and cuZFP (Lindstrom, 2014; cuZFP, 2020) have been leveraged by existing scientific applications, and cuSZp (Huang et al., 2023, 2024a) is an excellent high-throughput data compressor to be leveraged in performance-favored GPU data compression. Nevertheless, a significant research gap remains in this field: An open-source design for high-ratio scientific error-bounded lossy compression on GPUs is still missing. Most existing works focus on high throughput, relying on low-cost and low-quality data processing algorithms. cuSZ-I (Liu et al., 2024c), as the first and the only existing GPU-based high-ratio-and-quality scientific data compressor, integrates an NVIDIA-proprietary lossless encoding module to achieve a compression ratio higher than 32, yet still exhibits a much lower compression ratio than CPU-based compressors like SZ3 (Liang et al., 2023) and QoZ (Liang et al., 2023; Liu et al., 2024a). The most critical challenges of high-ratio scientific error-bounded lossy compression on GPU platforms are: 1) It is non-trivial to propose effective GPU-parallelization models and high-throughput implementations of high-ratio scientific data compression; 2) To optimize the GPU-based high-ratio scientific lossy compression pipelines, a thorough and in-depth benchmarking and analysis on the data decomposition and encoding modules on GPU platforms is required.
@@
In this paper, committed to optimizing high-ratio scientific error-bounded lossy compression on GPU platforms, we propose cuSZ-Hi, which is a general, open-source, and highly effective solution. From the decomposition-encoding perspective of error-bounded lossy compression, based on the prior success of interpolation-based error-bounded lossy compressors (Zhao et al., 2021; Liu et al., 2022; Liu et al., 2024a, c), we fine-craft a well-rounded GPU-parallelized interpolation-based data predictor, which is oriented toward maximizing the lossless compressibility of integer decomposition data. Furthermore, from systematic investigation, evaluation, and benchmarking on GPU-based numerical lossless encoders, we establish the best lossless pipelines for cuSZ-Hi, which fully leverages different aspects of redundancy information in the data, ensuring the high-magnitude compression ratio among all use cases. Our contributions are as follows:
@@
  • We perform a comprehensive evaluation of cuSZ-Hi on 6 real-world scientific datasets. cuSZ-Hi exhibits an outperforming compression ratio and quality compared to all existing state-of-the-art. cuSZ-Hi can achieve up to 249% compression ratio improvement over existing scientific lossy compressors under the same error bound and up to 215% compression ratio improvement under the same decompression data PSNR.
We organize the rest of this paper as follows: §2 discusses the related work of this paper. §3 formulates our research target and the background of interpolation-based scientific error-bounded lossy compression on GPUs. In §4, we propose the framework design and module composition of cuSZ-Hi. §5 demonstrates our detailed designs for optimizing cuSZ-Hi. In §6, we present the evaluations for cuSZ-Hi. §7 concludes this work and discusses future plans.
@@ 2.1. Scientific error-bounded lossy compression
Error-bounded lossy compression has been regarded as the best-fit strategy for scientific data size reduction. Existing error-bounded lossy compressors have proposed diverse technical solutions, and present different capabilities on use cases and metrics. Among the most representative works, SZ-family (Zhao et al., 2020b; Liang et al., 2023; Liu et al., 2022; Liu et al., 2024a) follows a hybrid design of spline interpolation (Zhao et al., 2021), Linear Regression, and Lorenzo extrapolation (Tao et al., 2017). ZFP (Lindstrom, 2014) and SPERR (Li et al., 2023) process data with transforms and specialized encoding schemes. TTHRESH (Ballester-Ripoll et al., 2019) leverages high-order singular value decomposition to effectively decorrelate the input data. To maximally reduce the compressed data size, there are various deep-learning-integrated frameworks for scientific error-bounded lossy compression (Liu et al., 2021a, 2023; Huang and Hoefler, 2022; Lu et al., 2021; Hayne et al., 2021; Liu et al., 2024b; Li et al., 2024). To optimize the compression throughput, compressors like SZx (Yu et al., 2022) and SZp (FZ-light) (Huang et al., 2024b) reduce their lossy data processing scheme to very-simplified offset-computation and bit-analysis.
@@ 2.2. Error-bounded lossy compression on GPUs
  • cuSZ: cuSZ (Tian et al., 2020, 2021; Liu et al., 2024c) was proposed in 2020, revised in 2021, and updated in 2024. The latest version of cuSZ merges the recently proposed cuSZ-I (Liu et al., 2024c) and the original Lorenzo-based cuSZ (Tian et al., 2020, 2021). It integrates 2 data predictors (Lorenzo predictor and interpolation predictor) and then applies Huffman encoding (and optionally, NVIDIA Bitcomp) as the lossless module.
  • cuSZp2: cuSZp2 (Huang et al., 2024a) is an end-to-end performance optimized scientific error-bounded lossy compressor on GPU platforms. It leverages 1D offset prediction and fixed-length encoding for high-speed scientific data compression.
@@
  • Others: cuSZx (Yu et al., 2022) features a monolithic design, delivering extremely high throughput at the cost of lower data quality and compression ratio. MGARD-GPU (Gong et al., 2023) is another GPU compressor practice based on the MGARD compression algorithm (Ainsworth et al., 2018).
@@ 3.1. Problem formulation of scientific error-bounded lossy compression
[…]
In Eq. 1, the upper bound of point-wise data error $\epsilon$ is named the error bound in the compression task. We further specify this problem in 2 aspects: 1) Regarding the GPU-based error-bounded lossy compression, the whole compression pipeline should be highly parallelized on the hardware, presenting high computing throughputs (often ten or hundreds of Gigabytes of data per second in modern platforms); 2) We observe that scientific error-bounded lossy compressors mostly follow a framework design of lossy decomposition plus lossless encoding (detailed in Section 4). The input floating-point data is first lossily decomposed into integer data with high compressibility, and the integers are then further losslessly encoded to achieve size reduction. Eq. 2 involves this framework scheme, formulating our research target. We would like to jointly optimize the integer decomposition process (Dcp) and the lossless encoding (Enc) process while preserving the required computing throughput ($\operatorname{Speed}(\operatorname{Dcp},\operatorname{Enc})$) with speed higher than a threshold $S_{t}$ (we initially set it to a comparable speed to cuSZ-I):
@@ 3.2. Interpolation-based error-bounded lossy compression on GPUs
In this subsection, we introduce the interpolation-based error-bounded lossy compression on GPUs (specifically cuSZ-I (Liu et al., 2024c)) as our research background, as cuSZ-Hi also integrates an interpolation-based data predictor. The core design of cuSZ-I is illustrated in Figure 1. To ensure parallelization, cuSZ-I partitions the input data into small chunks (e.g., $33\times 9\times 9$ for 3D data), each composed of 4 basic blocks (shown in Figure 1), and allocates each chunk to a thread block. Leveraging spline interpolations, a thread block predicts the allocated data chunk from a losslessly stored sparse anchor grid (in the stride of 8 on each dimension) and then quantizes the prediction errors to integers for the subsequent lossless size reduction process. The interpolations follow a hierarchical scheme, performing interpolation in different sequential steps, from high levels (with large interpolation strides) to low levels (with small interpolation strides) along all dimensions. The computations are fully parallelized on the threads within each interpolation level and dimension. After acquiring the quantized integer errors (also known as the quantization codes), cuSZ-I applies Huffman encoding and an optional proprietary NVIDIA Bitcomp (nvc, [n.d.]) lossless compression module to finalize the size reduction.
cuSZ-I (especially its Bitcomp-integrated version) has achieved state-of-the-art compression ratio and data quality in GPU-based scientific error-bounded lossy compression. In this paper, we will keep focusing our compressor design on interpolation and integer encoding techniques due to the unique potential of interpolation in proposing high-accuracy data predictions with relatively low computational costs. Furthermore, we will address several existing critical limitations of interpolation-based scientific data compression on GPUs, including but not limited to 1) suboptimal data-partition and anchor-sampling strategy; 2) outdated design of the interpolation process and its auto-optimization; 3) lack of an effective and open-source lossless encoding solution for quantized errors.
@@ 4. The Framework Design of cuSZ-Hi
The compression framework design of cuSZ-Hi is proposed in Figure 2, comparing it to other existing representative GPU-based scientific error-bounded lossy compressors. Although those compressors have diverse compression framework designs, they all share a uniform scheme, which is lossy data decomposition and lossless integer encoding, as shown in Figure 2. The first module of each framework decorrelates the input floating-point data, producing integers as intermediate results. Those integers have higher compressibility than the original input, contain reduced information but preserve the error bound, and are generated by different data processing techniques such as direct value quantization (in cuSZp2), Lorenzo extrapolation (in cuSZ), and spline interpolation (in cuSZ-I). Afterward, the lossless encoding modules in each compressor reduce the size of the intermediate integers without any information loss, ensuring that the final output serves as the error-bounded compression result. We detail the framework modules of cuSZ-Hi as follows and explain why they bring an optimized design regarding compression ratio and quality.
@@
  • Lossy data decomposition: cuSZ-Hi adopts spline-interpolation-based data prediction, which features an advanced design over the predictor in cuSZ-I (Liu et al., 2024c). Spline-interpolation-based data prediction has been proven to be the best-fit data decomposition technique for scientific error-bounded lossy compression because it can achieve both high data prediction accuracy and computation speed (Zhao et al., 2021; Liu et al., 2022; Liu et al., 2024a, c). In contrast, compressors with other lossy decomposition techniques, such as cuSZ, cuSZp2, and cuZFP, suffer from a significantly lower compression ratio than cuSZ-I. Based on the high-throughput parallelized implementation of spline interpolations proposed by (Liu et al., 2024c), cuSZ-Hi further tunes and optimizes the GPU-based spline interpolation design to maximize the compression ratio acquired.
@@
  • Lossless integer encoding: To maximize the compression ratio, a well-rounded lossless encoding pipeline is necessary, which should be able to leverage the data correlation information on both the token-level and bit-level. Unfortunately, among existing compressors, cuSZ only integrates the frequency-based Huffman encoding, and cuSZp2/cuZFP merely applies fixed-length bit-wise de-redundancy encodings. cuSZ-I managed to combine both Huffman encoding and de-redundancy encoding, but the NVIDIA Bitcomp module it leverages is a proprietary product with limited usability. In cuSZ-Hi, after synthetically investigating and benchmarking available lossless encoding solutions, cuSZ-Hi proposes new open-source lossless pipelines that jointly facilitate Huffman encoding and high-performance multi-stage lossless modules. Moreover, as indicated in Figure 2, cuSZ-Hi integrates 2 different lossless pipelines, which can be flexibly selected according to the user’s requirements, presenting a higher compression ratio (in cuSZ-Hi-CR mode) or throughput (in cuSZ-Hi-TP mode).